What is realtivsitic kinetic energy
1
Expert's answer
2020-12-10T11:10:46-0500

The relativistic kinetic energy is given by

"E_{kinetic} = \frac{mc^2}{\sqrt{1-\frac{v^2}{c^2}}} - mc^2"

We see that it is a growing function of "v^2" , is 0 for "v=0" and goes to "+\infty" when "v" approaches "c" . We can even show that for "v \ll c" we have the classical expression of kinetic energy:

"E \underset{v/c\to 0}{\sim} mc^2 (1+\frac{1}{2}\frac{v^2}{c^2}) - mc^2 = \frac{mv^2}{2}"

The relativistic kinetic energy have the same physical meaning: it is the energy of an object related to it's movement independently of a direction.